Solve the following equations:
2q + 6 = 0..............

Grade:12

2 Answers

Harshit Singh
3 years ago
Dear Student

First we have to subtract 6 to the both sides of the equation,
Then, we get,
= 2q + 6 - 6 = 0 - 6
= 2q = - 6
Now,
We have to divide both sides of the equation by 2,
Then, we get,
= 2q/2 = - 6/2
= q= - 3

Thanks
Shraddha Singh
2 years ago
Dear Student,
 
We can solve this question as follows :
2q + 6 = 0
2q = -6
q = -6/2
q = -3
 
Hope it helps!
Thank you ✌️